OpenStudy (anonymous):

Samantha made a triple batch of chocolate chip cookies. She used 7 and 5/8(mixed number) cups of flour. Before she made the cookies,she had 10 and 2/3(mixed number) cups of flour. How much flour does Samantha have left? Show your work.

OpenStudy (anonymous):

Subtract.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

10-7=? 3 5-2=? 3 8-3=? 5 answer 3 3/5

OpenStudy (anonymous):

Thank you for medal. :)

OpenStudy (anonymous):

That's not the answer

OpenStudy (anonymous):

What's answer choices?

OpenStudy (anonymous):

There id none but the answer is 3 and 1/24

OpenStudy (anonymous):

oh gosh i'm so sorry. I thought that be answer...... My mistake thats my first time I messed up. What grade you in??

OpenStudy (anonymous):

It's fine but im 7th

OpenStudy (anonymous):

Ah. I'm about to figure out how you get 3 1/24